IPP Progress Report for the week 2013.02.18 - 2013.02.24

(Up to IPP Progress Reports)

Eugene Magnier

Serge Chastel

  • Off on Monday (MLK Holiday)
  • IPP
    • ippdb02 wreckage
    • Heather's plots reimplemented/added to ippMonitor
    • Most of the files monitoring databases are up to date. I need to add some control of the status of the mysql servers on the server nodes. I'd like to know why they randomly crash...
  • IPP/MOPS
    • Wasted about a day trying to understand why the modified IPP didn't work anymore on Tuesday.
    • The modified IPP can now process detrended images and compute their difference. The mops products are now outputs of the processing chain
    • Ran a bunch of tests on exposures. Still tuning it though.

Heather Flewelling

  • implemented batchID in all stages of ipptopsps batches (these go into the database)
  • testing of those batches
  • science stuff with CA/ROTSE
  • gave serge info on my addstar plots, info how to inject images into database
  • isp database testing: new datastore works, however, discovered old database was disabled. we are 1.5 months behind on downloading isp data
  • czartools for ipptopsps is broken - trying to work around that by looking at psps side to see how we are doing.
  • pspswg meeting
  • sick one day

Mark Huber

  • sick thursday
  • czar usual misc LAP monitor+fixing over long weekend and tuesday
  • continuing to trace through relevant IPP bits for WS diffim photometry issues, notes continue to be added to MD_WSdiffim_photcheck_test20121004 -- continued through the PSF generation to understand why the PSF photometry is not well reproduced and a poor PSF is determined from a match+merged input list rather than just a matched source list.

Bill Sweeney

  • Spent most of the week finishing up the database changes and tools for IPP release tracking
  • Reviewed Thomas' UI for the PSI postage stamp interface. Made some suggestions
  • wrote up notes on how the postage stamp preview mode will be implemented and requirements for request and job tracking operations.
  • modified skycalResult table to include seeing measurements. Wrote a script to read the 1.3 million or so cmfs to extract the data and the columns in the db

Chris Waters

  • Tracked down bug in stacking to a bug in the EXT_FIT_ALL code I put into psphot. This was causing all sources to be fit inappropriately, leading to all stars being flagged as extended objects, leading to poor choices of PSFs for matching at the stack stage. Fixed on trunk.
  • Changed how WS diffs are queued in difftool. Instead of locking the database once and using a temp table to hold all run/input data, load that information into difftool memory and do multiple transactions that operate on a single run at a time. This allows the database to serve other requests.
  • Ran background continuity/background replacement code on M31. JPEGs are not automatically generated for background replacement, so I had to generate them manually. There are obvious issues on the scale of cells around large bright objects, likely due to the cell-level continuity not dealing with these steep gradients.
  • GPC1_M31_OTA_Offset_Study Looked into the reports that there are OTA level shifts in M31 images that have not had a background subtracted. I'm reasonably certain that this is due to problems in the dark model.